If a bullet leaves the muzzle of a rifle with a speed of 580. m/s and accelerates at a rate of 1.9 x 105 m/s2, how long is the barrel of the rifle? 


1
Expert's answer
2021-02-22T10:23:00-0500

Answer

By using

"v^2=u^2+2as"


initial velocity u=0m/s

By putting value

"(580)^2=(0)^2+2(1.9\times10^5)s\\s=\frac{580^2}{2(1.9\times10^5)}"


"s=0.88m"

So length of barrel is 0.88m